For the following problem, the slope and one point on a line is given. Use the point-slope form to find the equation of that line. (Write your answer in slope-intercept form.)
(−6,5), m = −43
Find y if the line through (5, 7) and (9, y) has a slope of 6.
Please help me im so confused :(

Answer by Alan3354(69443)

For the following problem, the slope and one point on a line is given. Use the point-slope form to find the equation of that line. (Write your answer in slope-intercept form.)
(-6,5), m = -43
-----
y - y1 = m*(x-x1) where m = slope and (x1,y1) is the point
---
y-5 = -43*(x+6) = -43x - 258
y = -43x - 253
======================================
Find y if the line through (5,7) and (9,y) has a slope of 6.
Slope = change in y/change in x
6 = (7-y)/(5-9) = (7-y)/(-4)
-24 = 7-y
-31 = -y
y = 31
----------
Don't be nervous, it's just arithmetic.